Show that if A is an n × n matrix, then AAT

and A + A

T

are symmetric.


1
Expert's answer
2021-05-12T03:48:14-0400

Let A be an "n\times n" matrix.


1. Let "B=AA^T"

  Take transpose on both sides-

   "B^T=(AA^T)^T\\

 B^T=(A^t)^T.A^T\\

 B^T=AA^T"


    As, "B^T=B=AA^T"


  So "AA^T" is symmetric.


2.Let "C=A+A^T"

   Take transpose on both sides-


 "C^T=(A+A^T)^T\\

 C^T=A^T+(A^T)^T\\

 C^T=A+A^T"


 hence "C=C^T=A+A^T"


So "A+A^T" is symmetric.